OpenStudy (anonymous):

What is 6.4*10^-8 mi/ day in feet per hour??

OpenStudy (anonymous):

\[\frac{6.4\times10^{-8}\text{ mi}}{1\text{ day}}\times\frac{5280\text{ ft}}{1\text{ mi}}\times\frac{1\text{ day}}{24\text{ hr}}\] With scientific notation, that's \[\frac{6.4\times10^{-8}\text{ mi}}{1\text{ day}}\times\frac{5.28\times10^3\text{ ft}}{1\text{ mi}}\times\frac{1\text{ day}}{2.4\times10^1\text{ hr}}=\frac{6.4\times5.28}{2.4}\times10^{-6}\] \[\frac{6.4\times5.28}{2.4}=14.08=1.408\times10^1\] \[(1.408\times10^1)\times10^{-6}=1.408\times10^{-5}\] Make any necessary adjustments if you're accounting for significant digits.